OpenStudy (anonymous):

Can someone help with with a question please? >Using Quadratic Formula and/or Substitution solve: ((t^2) -7) - 3((t^2) -7) + 2 = 0 (t^2) = T to the second power

OpenStudy (amistre64):

expand the problem and see what it gets you

OpenStudy (anonymous):

heres what i got so far. : Using the quadratic formula: (Substitution) U = ((t^2) - 7) ==> u^2 - 3u + 2 ==> Quadratic Formula ==> 3 +- sqrt{9+8 ==> Answer that i got was : 3 (plus or minus) sqrt{ 17} divided by 2

OpenStudy (anonymous):

Is this correct?

OpenStudy (amistre64):

not according to the question you posted; is the original typoed?

OpenStudy (anonymous):

hmm shoot yes it is, im so sorry. its : ((t^2) - 7)^2 - 3((t^2) -7) + 2 = 0

OpenStudy (amistre64):

ok..... your quadratic formula has issues to it; you seem to have forgotten to /2a

OpenStudy (anonymous):

Um at the end of my answer i stated "Divided by 2"

OpenStudy (amistre64):

i see it; its was lost in the formating lol

OpenStudy (amistre64):

sqrt(b^2 -4(a)(c))...youve got a +4(a)(c)

OpenStudy (amistre64):

3 sqrt[9-8] --- +- -------- 2 2

OpenStudy (anonymous):

Wow, i cant believe i made that mistake.

OpenStudy (amistre64):

your answers should end up as : t^2 -7 = 1 and t^2 -7 = 2

OpenStudy (amistre64):

you slow typing ? :)

OpenStudy (anonymous):

May i ask, after the 3+- sqrt { 9-8} = 3+- sqrt {1} ----------- 2 so ==> it'd be 3+ 1 \2 || 3-1 \2?

OpenStudy (anonymous):

im at work sorry lol

OpenStudy (amistre64):

lucky you; i aint got a job ...

OpenStudy (amistre64):

yes; as long as i see it correctly; you did that right

OpenStudy (anonymous):

:D thanks for your help. And oh on that note, ill be unemployed by the end of this month. :( Work place is closing shop. ----- Quick confirmation: final answer = final answers = 2 and 1 correct?

OpenStudy (amistre64):

the final answer is to solve for t....not u :) final answers should be: +- 2sqrt(2) and +-3

OpenStudy (anonymous):

ah... if i may, how did you arrive at that?

OpenStudy (amistre64):

u = 1 = t^2-7 u=2=t^2-7 8=t^2 9=t^2 +-2sqrt(2) +-3

OpenStudy (anonymous):

:D thanks man. i get it (mostly :P) this sites awesome :) i think i might post here for help some other time

OpenStudy (amistre64):

:) good luck